# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### v1

@@ Abstract
Understanding the dynamical origin of high–temperature superconductivity remains a central challenge of strongly correlated quantum matter. Conventional approaches to quantum criticality assume overdamped Markovian dissipation governed by Ohmic Landau damping. Here we show that infrared collective dynamics is instead generically controlled by the time–scale density of states (TDOS) of relaxation modes. Within the Martin–Siggia–Rose–Janssen–De Dominicis formalism, we derive an exact spectral representation of the collective susceptibility in terms of the TDOS. A finite infrared TDOS defines a new universality class of memory–dominated critical dynamics characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. This spectral reorganization produces a strong infrared amplification of the particle–particle channel, converting the marginal instabilities of BCS and Eliashberg theories into algebraic superconducting transitions. As a result, the transition temperature scales linearly with the infrared spectral weight of slow collective modes, naturally yielding superconducting domes and Uemura scaling without invoking bosonic glue or fine tuning. The same slow–mode reservoir governs anomalous normal–state dynamics, including long–time correlations and strange–metal behavior, providing a unified description of thermodynamic and dynamical phenomena in correlated superconductors. Our results establish dynamical spectral organization as a fundamental principle of quantum critical matter and identify memory–dominated criticality as a generic route to high–temperature superconductivity.
@@
A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11]. Such low-dimensional reductions necessarily generate a finite set of soft modes and therefore preclude the emergence of a dense continuum of relaxation time scales. A genuine slow-mode spectrum can only arise when the full high-dimensional collective dynamical state space is retained.
@@
In this work we show that infrared collective dynamics is generically controlled not by a single overdamped mode but by the time–scale density of states (TDOS) of correlated internal relaxation modes. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [12, 13, 14], we derive an exact spectral representation of the collective susceptibility in terms of the TDOS.
@@
We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics. While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$. This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a new universality class governed by self–organized slow modes rather than fast external baths.
@@
Crucially, we show that such flat relaxation mode spectra arise naturally in strongly correlated collective systems, including self–organized dynamical manifolds explored in our previous work [15, 16, 17]. The resulting long–memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a natural route to strong superconductivity without invoking finely tuned bosonic mediators.
@@
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation mode spectra determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 18, 19]. Charge, spin, and pairing channels reorganize simultaneously within the same slow–mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a common infrared spectral continuum.
@@
Our results establish the relaxation mode spectra as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified perspective on collective behavior across strongly correlated quantum matter [5, 7, 10].
@@
Strongly correlated electronic matter hosts many intertwined collective fluctuation channels, including charge, spin, current, pairing, and phase–coherence sectors. These degrees of freedom continuously exchange energy and information and cannot, in general, be reduced to a small set of independent order parameters.
To capture this collective complexity in a unified manner, we represent the correlated system as a high–dimensional interacting dynamical system whose instantaneous collective configuration is encoded in a state vector
$x(t)\in\mathbb{R}^{N},$ (1)
which evolves under nonlinear many–body feedback.
The coarse-grained dynamics may therefore be written as a stochastic field theory
$\dot{x}(t)=F(x(t))+\xi(t),$ (2)
where $F$ encodes correlated interactions and $\xi$ is Gaussian noise with
$\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=2D\,\delta_{ij}\delta(t-t^{\prime}).$ (3)
@@
Near correlated critical regimes the system does not soften a single collective mode but undergoes a reorganization of its full dynamical stability structure. Linearizing around a fluctuating metastable configuration $x^{*}$,
$\delta\dot{x}(t)=-A\,\delta x(t)+\xi(t),$ (4)
with
$A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}},$ (5)
yields a spectrum of collective relaxation modes
$Av_{i}=\lambda_{i}v_{i},$ (6)
whose retarded propagators are
$G_{i}(\omega)=\frac{1}{-i\omega+\lambda_{i}}.$ (7)
The infrared structure of correlated dynamics is therefore controlled by the distribution of relaxation rates $\lambda_{i}$. We characterize this spectrum by the time-scale density of states,
$\rho(\lambda)=\Big\langle\sum_{i}\delta(\lambda-\lambda_{i})\Big\rangle.$ (8)
In weakly interacting systems the TDOS vanishes as $\lambda\to 0$, reflecting a finite number of quasiparticle-like modes. By contrast, strongly correlated critical regimes generically generate an extensive accumulation of near-marginal directions in phase space,
@@
forming a slow-mode reservoir that governs infrared collective dynamics.
This accumulation is the dynamical signature of correlated criticality: the system approaches marginal stability along many collective directions rather than along a single order-parameter coordinate.
II.2 B. Exact spectral representation from MSRJD field theory
To compute collective response functions while retaining the full dynamical spectrum, we employ the Martin–Siggia–Rose–Janssen–De Dominicis formalism associated with Eq. (2),
$S[x,\tilde{x}]=\int dt\,\tilde{x}(t)\big[\dot{x}(t)-F(x(t))\big]-D\tilde{x}^{2}(t).$ (10)
We focus on a slow collective field $\phi(\mathbf{r},t)$ obtained by coarse graining the full state vector and regard the remaining fast correlated degrees of freedom as internal relaxation modes.
The reorganized stability spectrum may be represented exactly as a continuum of Ornstein–Uhlenbeck modes $X_{\lambda}$ labeled by their decay rates,
@@
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda}.$ (12)
Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action becomes
@@
$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi]$
@@
Since the reservoir variables enter quadratically, they can be integrated out exactly. In Fourier space this yields the effective collective action
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots,$ (14)
with retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (15)
@@
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\frac{\omega}{\lambda^{2}+\omega^{2}},$ (16)
gives
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (17)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{1}{\lambda^{2}+\omega^{2}}.$ (18)
Identifying the TDOS as
$\rho(\lambda)\equiv g(\lambda)^{2},$ (19)
the collective inverse susceptibility becomes
@@
Equation (20) constitutes an exact spectral representation of collective infrared dynamics in terms of the dynamical stability spectrum. The complete field-theoretic derivation of the collective susceptibility kernel from the MSRJD formalism, including the explicit Gaussian integration over the relaxation-mode reservoir and the resulting self-energy structure, is presented in Appendix A.
@@
Here $\omega$ denotes the frequency conjugate to the collective relaxation dynamics obtained after integrating out the slow–mode reservoir in the MSRJD formulation, rather than a microscopic quasiparticle energy scale.
@@
in direct analogy with the conventional Hertz–Millis representation. The TDOS scaling immediately implies the universal identification $s=\alpha$.
@@
In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the presence of an extensive reservoir of slow collective relaxation modes. Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$. Infrared universality classes are thus classified directly by the low–energy organization of relaxation modes encoded in the TDOS.
@@
corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, which is derived explicitly in Appendix B.
This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory. Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow internal modes.
The TDOS thus provides a unified organizing principle for infrared critical dynamics, with different universality classes emerging from the scaling of the dynamical stability spectrum itself.
@@
The emergence of a slow–mode reservoir qualitatively reorganizes the infrared structure of the Cooper pairing channel. In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal. Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
@@
The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining. Physically, overdamped collective modes rapidly dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
@@
In the memory–dominated regime derived in Sec. II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$. After integrating out the relaxation reservoir within the MSRJD framework, the retarded collective susceptibility acquires the exact spectral form
@@
Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct. In Eliashberg theory the singularity reflects instantaneous overdamped dissipation arising from coupling to a fast electronic bath [20, 21]. Here, by contrast, the same infrared scaling is a consequence of self–organized spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
@@
The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one. This qualitative contrast with Eliashberg theory is illustrated in Fig. 2. A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix C.
Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and dynamically amplify intrinsic electronic pairing tendencies.
Superconductivity in strongly correlated systems thus emerges from infrared spectral self–organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
@@
In this section we show explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, yielding power–law growth of the effective pairing interaction and algebraic scaling of the transition temperature.
@@
More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated by electronic interactions, independent of long–wavelength collective mediators. This intrinsic scale provides the microscopic seed for superconductivity.
@@
with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies this preexisting pairing tendency.
@@
Because single–particle propagators admit a relaxational spectral representation controlled by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
@@
This nonanalytic enhancement reflects the accumulation of collective memory across all time scales. Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel is dominated by long–time memory effects generated by the slow–mode TDOS.
IV.3 C. Algebraic superconducting instability and universality
@@
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
@@
Because the TDOS exponent controls infrared dynamics independently of microscopic details, the pairing amplification mechanism is universal across correlated systems exhibiting flat relaxation spectra. The slow–mode reservoir therefore acts as a universal dynamical amplifier, converting electronic correlation energy into enhanced superconducting transition temperatures without invoking specific bosonic glue.
@@
The linearized transition scale derived in Eq. (45) shows that the superconducting temperature is controlled by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. Near correlated quantum criticality, $\rho_{0}$ is not constant but is governed by universal scaling.
Let $p$ denote a tuning parameter (e.g., doping or pressure) and $p_{c}$ the critical point. The mass term of the collective kernel obeys
$r(p)\sim|p-p_{c}|,$ (46)
so that the corresponding infrared cutoff scale takes the universal form
$\Lambda(p)\sim r^{z\nu}\sim|p-p_{c}|^{z\nu},$ (47)
with $z$ the dynamical exponent and $\nu$ the correlation-length exponent. Away from criticality the slow–mode reservoir is therefore truncated in the infrared, yielding an effective low-energy TDOS weight
$\rho_{0}(p)\sim\int_{0}^{\Lambda(p)}d\lambda\,\rho(\lambda).$ (48)
For a memory–dominated regime in which the TDOS remains finite at the origin, $\rho(\lambda\to 0)=\rho_{*}$, one obtains parametrically
$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{p_{0}}\right),$ (49)
where $f$ is a universal decreasing scaling function and $p_{0}$ denotes a nonuniversal crossover scale.
Substituting Eq. (49) into Eq. (45) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{x_{0}}\right),$ (50)
which is maximal near criticality and suppressed on both sides, producing a superconducting dome without fine tuning.
The dome structure is thus a direct consequence of the competition between intrinsic short–range electronic pairing and the finite infrared extent of the slow–mode reservoir away from criticality. Importantly, no competing order parameters or additional microscopic suppression mechanisms are required: suppression of $T_{c}$ arises purely from the truncation of long–time memory encoded in the TDOS. In this interpretation, optimal doping corresponds to maximal infrared spectral reorganization rather than to the peak of any specific bosonic fluctuation.
@@
While the microscopic pairing seed is set by the intrinsic electronic scale $E_{\rm pair}$, the establishment of superconductivity requires global phase coherence of the Cooper collective field. In systems with low superfluid density, the transition may be controlled by phase stiffness rather than by gap formation.
The effective phase–only action for the Cooper field reads
@@
where $\rho_{s}$ is the superfluid stiffness. The phase–ordering temperature therefore scales as
@@
Within the slow–mode reservoir framework, the stiffness is controlled by the same infrared spectral weight that enhances pairing. Integrating out amplitude fluctuations yields the scaling
@@
reflecting the collective participation of slow modes in coherent transport. Combining Eqs. (52) and (53) gives

### v2

# arXiv:2602.22626v2[cond-mat.str-el] 02 Mar 2026

@@ Abstract
[…]
@@
A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11]. Such descriptions implicitly assume that the infrared behavior can be captured by a finite set of soft bosonic modes with local-in-time damping. However, strongly correlated quantum matter is fundamentally governed by Liouvillian evolution of the density matrix, where the full spectrum of relaxation eigenvalues determines long-time dynamics [12, 13, 14]. In more general, criticality need not correspond to softening of a single excitation mode but may instead involve reorganization of the entire relaxation-rate spectrum.
Rather than reducing the dynamics to a few collective coordinates, one must consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes. Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality. The relevant organizing structure of the infrared theory is then not a small set of overdamped bosons, but the time–scale density of states (TDOS) of collective relaxation modes.
@@
In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
@@
We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics. While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$. This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a distinct universality class governed by the structure of the relaxation spectrum.
@@
Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [18, 19, 20]. The resulting long-memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a route to strong superconductivity without invoking finely tuned bosonic mediators.
@@
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 21, 22]. Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
@@
Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 10].
@@
Strongly correlated electronic matter is intrinsically a quantum many-body system. At long times and coarse-grained scales, its dynamics is governed not by isolated quasiparticles, but by collective relaxation channels generated by the full quantum dynamical evolution of the density matrix [12, 13, 14],
$\frac{d\hat{\varrho}}{dt}=\mathcal{L}\hat{\varrho}$, (1)
where $\mathcal{L}$ denotes the Liouvillian generator of the effective (open or coarse-grained) dynamics.
The eigenmodes of $\mathcal{L}$,
$\mathcal{L}R_{n}=\lambda_{n}R_{n},$ (2)
define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as
$\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0.$ (3)
Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
Observable collective responses—charge, spin, pairing, current, and phase-coherence fluctuations—arise from projections of the full relaxation spectrum onto experimentally accessible operators. The resulting low-energy theory is thus controlled by the statistical structure of relaxation rates. We define the TDOS as the infrared distribution of Liouvillian relaxation eigenvalues,
$\rho(\lambda)=\sum_{n}\delta\!\left(\lambda+\mathrm{Re}\,\lambda_{n}\right),$ (4)
or its continuum limit.
The spectral structure introduced here should therefore be understood at the level of the full dynamical generator rather than solely in terms of Hamiltonian energy eigenmodes. Energy spectra characterize the conservative limit in which all Liouvillian eigenvalues lie on the imaginary axis, whereas generic interacting or coarse-grained systems develop complex spectra with nonzero decay rates. A detailed discussion of this relaxation-rate (Liouvillian) formulation and its connection to memory kernels is provided in Appendix A.
@@
The infrared structure of correlated matter is encoded in the spectrum of collective relaxation rates. Rather than softening a single order parameter, correlated critical regimes reorganize the entire decay spectrum, producing a dense accumulation of near-marginal modes.
We characterize this structure by introducing the TDOS, which measures the density of collective decay channels at relaxation rate $\lambda$. This object plays a role analogous to the energy density of states, but in relaxation-rate space. It determines how many collective modes decay on a given time scale.
In weakly interacting systems, the number of slow modes remains finite, and the TDOS vanishes as $\lambda\to 0$, reflecting the presence of well-defined quasiparticles and isolated collective excitations.
By contrast, in strongly correlated critical regimes, the relaxation spectrum generically develops an extensive accumulation of near-marginal decay channels,
@@
forming a slow-mode reservoir that dominates infrared collective dynamics.
This accumulation signifies a dynamical approach to marginal stability along a macroscopic manifold of collective directions, rather than along a single order-parameter coordinate.
II.2 B. Functional representation of the relaxation spectrum
To compute collective response functions, we employ a functional representation of the dissipative dynamics. The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework [23, 24]. Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
Let $\phi(\mathbf{r},t)$ denote a coarse–grained collective bosonic field representing charge, spin, pairing, or phase fluctuations. Projecting the full quantum dynamics onto this operator sector yields an exact decomposition in terms of relaxation eigenmodes.
Because the dissipative spectrum consists of decay rates $\lambda\geq 0$, the projected dynamics may be represented as a continuum of relaxational modes labeled by their decay scales,
@@
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda},$ (7)
where $X_{\lambda}$ represents a decay mode with relaxation rate $\lambda$, and $g(\lambda)$ encodes its overlap with $\phi$.
Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action reads
@@
$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}\big[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi\big]$
@@
Because the reservoir fields $X_{\lambda}$ enter quadratically, they can be integrated out exactly. Performing the Gaussian functional integration generates a nonlocal-in-time memory kernel in the effective equation of motion,
$\partial_{t}\phi(t)=-(r+c\nabla^{2})\phi(t)-\int_{0}^{t}dt^{\prime}\,K(t-t^{\prime})\,\phi(t^{\prime})+\eta_{\rm eff}(t),$ (9)
where the memory kernel is given exactly by
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (10)
with the time-scale density of states
$\rho(\lambda)\equiv g(\lambda)^{2}.$ (11)
Equation (10) makes explicit that the slow-mode reservoir is entirely controlled by the distribution of relaxation rates. The long-time structure of $K(t)$ therefore encodes the infrared organization of decay modes in the interacting system.
Fourier transforming the kernel yields the retarded self-energy,
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
so that the effective action becomes
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
@@
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
The collective inverse susceptibility therefore assumes the exact spectral representation
@@
Equation (16) provides an exact infrared response fully controlled by the relaxation-rate density. The slow-mode reservoir thus appears as a memory-dominated fixed point, whose long-time kernel structure governs the dynamical amplification mechanism discussed below.
The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
@@
Here $\omega$ denotes the collective relaxation frequency arising from the projected Liouvillian dynamics, rather than a microscopic quasiparticle excitation energy.
@@
in direct analogy with the conventional Hertz–Millis representation [3, 4]. The TDOS scaling immediately implies the universal identification $s=\alpha$.
@@
In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the organization of the underlying relaxation spectrum. Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$. Infrared universality classes are thus classified directly by the low–energy structure of the relaxation-rate density.
@@
corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory. Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow decay modes.
@@
The emergence of a slow–mode reservoir reorganizes the infrared structure of the Cooper pairing channel. In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal. Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
@@
The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining. Physically, overdamped collective modes dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
@@
In the memory–dominated regime derived in Sec. II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$. After integrating out the relaxation reservoir, the retarded collective susceptibility takes the exact spectral form
@@
Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct. In Eliashberg theory the singularity reflects overdamped dissipation arising from coupling to a fast electronic bath [25, 26]. Here, by contrast, the same infrared scaling results from spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
@@
The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one. This qualitative contrast with Eliashberg theory is illustrated in Fig. 2. A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix D.
Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and amplify intrinsic electronic pairing tendencies.
Superconductivity in strongly correlated systems thus emerges from infrared spectral organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
@@
In this section we demonstrate explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, ultimately yielding algebraic scaling of the transition temperature and robust phase coherence.
@@
More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated purely by electronic interactions, independent of long–wavelength collective mediators. This intrinsic scale provides the microscopic seed for superconductivity.
@@
with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
@@
Because single–particle propagators admit a relaxational spectral representation governed by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
@@
This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales. Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel becomes relevant due to the flat TDOS. The slow–mode reservoir therefore reorganizes the pairing problem into a memory–dominated critical regime.
IV.3 C. Algebraic Instability and Transition Scale
@@
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix E.
@@
It is important to emphasize that superconductivity requires not only pair formation but also global phase coherence. The transition temperature is therefore ultimately bounded by the smaller of the pairing amplitude and the superfluid phase stiffness,
$k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right),$ (42)
with $\rho_{s}\sim n_{s}/m^{*}$. The detailed scaling of the phase stiffness within the slow–mode framework is discussed in Sec. V.
The key result of the present section is that the pairing scale itself is controlled by a single infrared parameter, the zero–frequency TDOS $\rho_{0}$. The slow–mode reservoir therefore acts as a universal dynamical amplifier: it converts intrinsic electronic correlation energy into an algebraically enhanced instability of the Cooper channel, independent of any specific bosonic mediator.
In this sense, the TDOS–controlled fixed point defines a dynamical stability class distinct from both weak–coupling BCS superconductors and conventional Eliashberg scenarios. The emergence of superconductivity is governed not by the exchange of a particular glue, but by infrared spectral organization of collective relaxation modes.
@@
The linearized transition scale derived in Eq. (41) shows that superconductivity is governed by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical fixed point of the relaxation spectrum.
In this phase, feedback between electronic correlations and collective relaxation continuously regenerates an extensive near–marginal sector, leading to a scale–free infrared TDOS,
$\rho(\lambda\to 0)=\rho_{*}>0,$ (43)
without the need for fine tuning of microscopic parameters. The flat infrared spectrum therefore characterizes a dynamical phase rather than a measure–zero critical point.
External control parameters such as doping or pressure primarily modulate the effective infrared cutoff scale of this reservoir. Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared cutoff is minimal. The collective mass term behaves as
$r(p)\sim|p-p_{\star}|,$ (44)
so that the associated cutoff obeys
$\Lambda(p)\sim r^{z\nu}\sim|p-p_{\star}|^{z\nu}.$ (45)
Away from $p_{\star}$, the slow–mode reservoir remains present but is truncated in the deep infrared, yielding
$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (46)
with $f$ a decreasing crossover function.
Substituting into Eq. (41) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (47)
which naturally produces a superconducting dome. Importantly, the suppression of $T_{c}$ away from $p_{\star}$ does not arise from the disappearance of the dynamical fixed point, but from the progressive truncation of its infrared extent.
In this interpretation, optimal doping does not correspond to a conventional symmetry–breaking quantum critical point. Rather, it marks the regime in which the self–organized infrared spectral phase is most fully exposed.
@@
While Sec. IV established that the pairing scale is algebraically enhanced by the infrared TDOS, the observable transition temperature is ultimately limited by global phase coherence. In systems with reduced superfluid density, the ordering temperature may be controlled by phase stiffness rather than by gap formation.
The long-wavelength phase dynamics of the Cooper field is described by the effective action
@@
where $\rho_{s}$ denotes the superfluid stiffness. The corresponding phase-ordering temperature scales as
@@
Within the slow–mode reservoir framework, the same infrared spectral weight that enhances pairing also governs coherent charge transport. The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore proportional to the density of coherent paired carriers that remain dynamically correlated over long time scales.
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$ [Eq. (39)], the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
@@
up to nonuniversal numerical factors.
Both pairing amplitude and phase rigidity are thus controlled by a single infrared parameter that quantifies the extent of the slow–mode reservoir.
Combining Eqs. (49) and (50) gives